A Company is considering leasing a new equipment. The lease lasts for 5 years. The lease calls for 5 payments of $41,000 per year with the first payment occurring immediately. The equipment would cost $192,000 to buy and would be straight-line depreciated to a zero salvage value over 5 years. The actual salvage value is negligible because of technological obsolescence. The firm can borrow at a rate of 6%. The corporate tax rate is 25%. What is the after-tax cash flow from leasing relative to the after-tax cash flow from purchasing in year 2?

$32,500

-$40,350

-$41,000

$27,500

-$42,750
Answer #1

-$40,350

Lease Option:
After tax lease rental in year 2 = Before Tax Lease rental*(1- Tax Rate) = -41000*(1-0.25) = $ -30,750
Lost Depreciation tax Shield = Depreciation*Tax rate = -38400*25% = $   -9,600
After tax cash flow in year 2 $ -40,350
Working:
Depreciation = (Cost - Salvage Value)/Useful life
= (192000-0)/5
= $ 38,400